For several days the temperature at the bottom of a pond is 4 degree Celsius and the air temperature just above the top surface of the pond is -1 Degree Celsius. Depth of the pond is L . The thermal conductivity of ice is 3 times that of water. Thickness of frozen layer of the ice formed must be?

Grade:10

1 Answers

Eshan
5 years ago
The rate of heat flowing from the bottom of the lake to the ice-water interface must be equal to that from the interface to the atmophere.

\implies \dfrac{(KA(4-0))}{L-x}=\dfrac{(3K)A(0-(-1))}{x}
\implies 4x=3(L-x)
\implies x=\dfrac{3L}{7}